OpenStudy (anonymous):

I need to make a presentation on MEAN in stats. Can someone gimme any ideas to make it interesting and informative??

OpenStudy (kira_yamato):

\[\bar{x} = \frac{\sum_{n=1}^{N}fx}{\sum_{n=1}^{N}f}\]

OpenStudy (anonymous):

hmmm i dont get it

OpenStudy (anonymous):

Mean is the average, it is the number that is obtained when all numbers are added together than divided by how many different numbers were added together. example Numbers 5,3,10,4,7,8,15 Total: 52 Mean = 52/7 = 7.43 If I was presenting this to a class I would ask people from the class to nominate different numbers write them up on the board and do the working to show the mean, which is simply the average. Hope it helps.